### v1

# Exotic Doublon-Holon Pairing State in Photodoped Mott Insulators

Abstract
We demonstrate the existence of a unique pairing state in photodoped Mott insulators on ladder geometries, characterized by quasi-long-ranged doublon-holon correlations, using the density matrix renormalization group method. This phase exhibits doublon-holon pairing correlations with opposite signs along the rung and chain directions, reminiscent of $d$-wave pairing in chemically doped ladder systems. By constructing the phase diagram, we reveal that the doublon-holon pairing state emerges between the spin-singlet phase and the charge-density-wave/$\eta$-pairing phase. Our study suggests that the interplay of charge, spin, and $\eta$-spin degrees of freedom can give rise to exotic quantum many-body states in photodoped Mott insulators.
External field driving is a useful tool for generating and controlling intriguing out-of-equilibrium phenomena in correlated electron systems [1, 2, 3, 4, 5, 6]. Photoinduced phase transitions [7, 8, 9, 10, 11, 12, 13], Floquet engineering [14, 15, 16, 17, 18], and high-harmonic generation [19, 20, 21, 22, 23, 24, 25, 26] have been extensively studied to date. In insulating materials, optical excitation across the band gap generates electron and hole carriers. When the recombination time is sufficiently long, the excited state potentially settles into a quasi-steady photodoped state, nearly conserving the number of carriers [27, 28]. Photoexcitation in correlation-driven Mott insulators (MIs) creates two types of carriers, doublons (doubly occupied sites) and holons (empty sites), see Fig. 1(b) [29, 30, 6, 31, 32, 33, 34], resulting in a unique photodoped state. In a large-gap MI, the carriers are expected to have long lifetimes, leading to a quasi-steady state where doublons and holons coexist [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
Figure 1: (a) Phase diagram of the photodoped ladder-type MI as a function of the doublon density $n_{d}$ and the intersite Coulomb interaction $V$ with the blue, red, green, and purple points representing the predominance of the spin, $\eta$-pairing, CDW, and doublon-holon (DH) pairing correlations, respectively. At the orange point ($n_{d}=0.5$ and $V=0.2t_{\parallel}$), the $\eta$-pairing and CDW states are degenerate. Parameters: $t_{\perp}=t_{\parallel}$, $J_{\perp}=J_{\parallel}=0.4t_{\parallel}$, and $V=V_{\parallel}=V_{\perp}$. (b) DH pairing state in the photodoped ladder with doublons and holons generated by light. The DH pairing correlation is characterized by an operator that maps spin singlets on nearest-neighbor sites to DH $\eta$-triplets.
Photodoping can be a trigger for creating exotic many-body states that are inaccessible in equilibrium, such as the $\eta$-pairing state [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The photodoped states have been investigated in detail in the one-dimensional (1D) Hubbard chain, where the phase diagram comprises the spin density wave (SDW) phase at low doublon density $n_{d}$ and the charge density wave (CDW) and $\eta$-pairing phases from intermediate to high $n_{d}$ [43, 44]. Although even richer physical properties are anticipated in the two-dimensional (2D) systems because of the increased geometric degrees of freedom and the lack of spin-charge-$\eta$ separation [44], these properties of photodoped 2D MIs are yet to be understood.
Ladder geometry can serve as a bridge between 1D and 2D systems and is more manageable for numerical approaches. For example, density matrix renormalization group (DMRG) calculations, which enable us to obtain numerically precise low-dimensional wave functions, have shown that chemically hole-doped ladders exhibit pair correlations corresponding to $d$-wave superconductivity (SC) in a 2D square lattice [57, 58, 59, 60, 61]. These studies are an important stepping stone to research on $d$-wave SC in the Hubbard and $t$-$J$ models on multileg cylinders, mimicking the 2D square lattice of cuprate superconductors [62, 63, 64, 65, 66, 67]. Furthermore, the pairing mechanism in ladder and bilayer systems [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79] has attracted significant attention associated with the recent discovery of high-temperature SC in bilayer nickelates [80, 81, 82, 83]. Because a strong interchain spin coupling favors the formation of spin-singlet rungs, hole doping to the ladder system leads to the formation of interchain hole-hole pairs to minimize the disruption of spin-singlet bonds, resulting in the development of SC correlations [68]. Realization of such a pairing mechanism has been recently proposed also with cold atoms in optical lattices [84]. From these considerations, we arrive at the crucial question: what are the equivalent paired states in photodoped MIs?
In this Letter, we show that photodoped MIs in the ladder geometry and quasi-stationary situation exhibit an exotic pairing state composed of doublons and holons (Fig. 1). This pairing is characterized by the quasi-long-range correlation of doublon-holon (DH) pairs on the spin background shown in Fig. 1(b). We find opposite phases for the rung and chain pairing, analogous to signatures of $d$-wave pairing in chemically doped ladders. We reveal the roles of interchain coupling and propose an effective minimal model that can capture the pairing properties.
Model and Methods—To model MIs in the presence of nonlocal interactions, we start with the extended Hubbard Hamiltonian in a ladder geometry
^Hub=subscript^$\hat{\mathcal{H}}_{\mathrm{Hub}}=$ $-t_{\parallel}\sum_{j,\alpha,\sigma}\left(\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j+1,\alpha;\sigma}+\mathrm{H.c.}\right)$
$-t_{\perp}\sum_{j,\sigma}\left(\hat{c}^{\dagger}_{j,0;\sigma}\hat{c}_{j,1;\sigma}+\mathrm{H.c.}\right)+U\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha;\uparrow}\hat{n}^{\prime}_{j,\alpha;\downarrow}$
$+V_{\parallel}\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha}\hat{n}^{\prime}_{j+1,\alpha}+V_{\perp}\sum_{j}\hat{n}^{\prime}_{j,0}\hat{n}^{\prime}_{j,1}.$ (1)
Here, $\hat{c}^{\dagger}_{j,\alpha;\sigma}$ ($\hat{c}_{j,\alpha;\sigma}$) is the creation (annihilation) operator for a fermion with spin $\sigma=\uparrow$, $\downarrow$ at site $j$ on chain $\alpha=0,1$. $\hat{n}_{j,\alpha;\sigma}=\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j,\alpha;\sigma}$, $\hat{n}^{\prime}_{j,\alpha;\sigma}=\hat{n}_{j,\alpha;\sigma}-1/2$, and $\hat{n}^{\prime}_{j,\alpha}=\hat{n}^{\prime}_{j,\alpha;\uparrow}+\hat{n}^{\prime}_{j,\alpha;\downarrow}$. $t_{\parallel}$ and $t_{\perp}$ are the hopping integrals along the chain and rung directions, respectively. $U>0$ is the on-site Coulomb interaction. $V_{\parallel}>0$ and $V_{\perp}>0$ are the nearest-neighbor Coulomb interactions along the chain and rung directions, respectively.
In photodoped MIs, doublons and holons are generated by an optical excitation, as shown in Fig. 1(b). When $U\gg t_{\parallel},t_{\perp}$, the recombination time is sufficiently long [35, 36, 37, 38, 39, 41, 40, 42] and the system can be considered to be in a pseudoequilibrium state where doublons and holons are conserved [43, 44]. Assuming $U\gg t_{\parallel},t_{\perp}$, we introduce an effective model derived by the Schrieffer-Wolff transformation that excludes the hopping terms changing the number of doublons and holons [39, 41, 43, 85]. The Hamiltonian of the effective $t$-$J$-$V$ model is given by
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}^{(0)}_{t}+\hat{\mathcal{H}}^{(s)}_{J}+\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}.$ (2)
Here, ^$\hat{\mathcal{H}}^{(0)}_{t}=\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}+\hat{\mathcal{H}}^{(0)}_{t_{\perp}}$ denotes the doublon-number-conserving hopping, see Supplemental Material for details [86], and ^$\hat{\mathcal{H}}_{V}=\hat{\mathcal{H}}_{V_{\parallel}}+\hat{\mathcal{H}}_{V_{\perp}}$ represents the nearest-neighbor Coulomb interactions $V_{\parallel}$ and $V_{\perp}$ in Eq. (1). ^$\hat{\mathcal{H}}^{(s)}_{J}=\hat{\mathcal{H}}^{(s)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}$ describes the Heisenberg-type spin-exchange interaction, while ^$\hat{\mathcal{H}}^{(\eta)}_{J}=\hat{\mathcal{H}}^{(\eta)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}$ represents the $\eta$-spin interaction for doublons and holons [43, 85]. These nearest-neighbor interactions between sites $\bm{i}$ and $\bm{j}$ are given by
^$\hat{\mathcal{H}}^{(s)}_{J_{\bm{i}\bm{j}}}=J_{\bm{i}\bm{j}}\left(\hat{\bm{s}}_{\bm{i}}\cdot\hat{\bm{s}}_{\bm{j}}-\frac{1}{4}\delta_{1,\hat{n}_{\bm{i}}\hat{n}_{\bm{j}}}\right),$ (3)
^$\hat{\mathcal{H}}^{(\eta)}_{J_{\bm{i}\bm{j}}}=-J_{\bm{i}\bm{j}}\left(\hat{\bm{\eta}}_{\bm{i}}\cdot\hat{\bm{\eta}}_{\bm{j}}-\frac{1}{4}\left(1-\delta_{1,\hat{n}_{\bm{i}}}\right)\left(1-\delta_{1,\hat{n}_{\bm{j}}}\right)\right),$ (4)
where $\bm{j}$ denotes $(j,\alpha)$. The spin operator is defined as $\hat{\bm{s}}_{\bm{j}}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{j,\alpha;\sigma}\bm{\sigma}_{\sigma,\sigma^{\prime}}\hat{c}_{j,\alpha;\sigma^{\prime}}/2$, where $\bm{\sigma}$ is the vector of Pauli matrices. The $\eta$-spin operator $\hat{\bm{\eta}}_{\bm{j}}$ is given by $\hat{\eta}^{+}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}^{\dagger}_{j,\alpha;\downarrow}\hat{c}^{\dagger}_{j,\alpha;\uparrow}$, $\hat{\eta}^{-}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$, and $\hat{\eta}^{z}_{\bm{j}}=\left(\hat{n}_{j,\alpha}-1\right)/2$ [87]. The explicit Hamiltonian for the ladder system is presented in the Supplemental Material [86]. For the hopping integral $t_{\bm{i}\bm{j}}$, the coupling constant is characterized by $J_{\bm{i}\bm{j}}=4t^{2}_{\bm{i}\bm{j}}/U>0$ (at $V_{\bm{i}\bm{j}}=0$). The $V$ term takes the form $V_{\bm{i}\bm{j}}\hat{n}^{\prime}_{\bm{i}}\hat{n}^{\prime}_{\bm{j}}=4V_{\bm{i}\bm{j}}\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}$, thereby introducing Ising-type anisotropy of $\eta$-spin. In contrast to the chemically doped $t$$J$ model, $\hat{\mathcal{H}}$ includes the interaction between doublons and holons described by $\eta$-spin, which enables the formation of $\eta$-pairing states in photodoped MIs [51, 43, 44].
We describe the photodoped ladder as the lowest-energy state of $\hat{\mathcal{H}}$ with a fixed doublon number. We obtain the state using the DMRG method [88, 89, 90]. Unless otherwise specified, we use $t_{\parallel}$ as the unit of energy and set the chain length to $L=160$ with open boundary conditions. The total number of sites is $2L$, and the numbers of up- and down-spin fermions are set to $N_{\uparrow}=N_{\downarrow}=L$ (i.e., half-filling without spin polarization), while the state is configured as a function of the doublon number. The bond dimension is set to $m=10000$ with the maximum truncation error typically being on the order of $1.0\times 10^{-7}$.
Numerical Results—To investigate which nonequilibrium phases can emerge in photodoped ladders, we study the behavior of correlation functions $C(r)=\braket{\hat{O}^{\dagger}_{j_{0}+r}\hat{O}_{j_{0}}}$ between operators $\hat{O}_{j}$ that are $r$ sites apart. We calculate correlations with respect to the site $j_{0}=L/4+1$ to minimize open boundary effects. The dominant nonequilibrium phase is determined by identifying the correlation function that exhibits the slowest decay with spatial distance $r$. Because the correlations between the single-site operators are symmetric on both chains, we consider the $\alpha=0$ chain index in the operators. We use $\hat{O}_{j}=\hat{n}_{j,\alpha;\uparrow}-\hat{n}_{j,\alpha;\downarrow}$ for the spin correlation, $\hat{O}_{j}=\hat{n}_{j,\alpha}-1$ for the charge correlation, and $\hat{O}_{j}=\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$ for the on-site pairing correlation.
Figure 2: (a) Log-log plot and (b) linear scale plot of the correlation functions for $n_{d}=0.2$, $t_{\perp}=t_{\parallel}$, $J_{\perp}=J_{\parallel}=0.4t_{\parallel}$, and $V=V_{\perp}=V_{\parallel}=0.2t_{\parallel}$. r-r (purple circles) and r-c (orange squares) represent the rung-rung and rung-chain DH pairing correlations, respectively. The on-site pairing correlations are multiplied by two for comparison on the same scale.
The central result of this Letter is the finding of the DH pairing phase in a finite region of doublon density $n_{d}=\sum_{j,\alpha}\braket{\hat{n}_{j,\alpha;\uparrow}\hat{n}_{j,\alpha;\downarrow}}/(2L)$ and at sufficiently strong nonlocal repulsion $V$, shown in Fig. 1(a). This DH pairing state is characterized by the correlation function $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ with $\hat{\Delta}^{\mathrm{r}}_{j}=\sum_{\alpha,\sigma}(-1)^{\alpha}\left(1-\hat{n}_{j,\bar{\alpha};\bar{\sigma}}\right)\hat{c}^{\dagger}_{j,\bar{\alpha};\sigma}\hat{c}_{j,\alpha;\sigma}\hat{n}_{j,\alpha;\bar{\sigma}}/2$, where $\bar{\alpha}=1$ $(0)$ for $\alpha=0$ $(1)$, and $\bar{\sigma}=\downarrow$ ()($\uparrow$)( ↑ ) for $\sigma=\uparrow$ ()($\downarrow$)( ↓ ). The operator $\hat{\Delta}^{\mathrm{r}\dagger}_{j}$ creates a rung $\eta$-triplet with a doublon and a holon on the rung spin-singlet background, as shown in Fig. 1(b) [91]. Here, the $\eta$-triplet states ($\eta=1$) with $\eta^{z}=-1,0$, and $+1$ correspond to holon-holon (HH), DH, and doublon-doublon (DD) pair states, respectively [86]. Considering ^$\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}$ on a single rung, we can see that the formation of the rung DH pair in the $\eta$-triplet state is energetically favorable in the presence of the nearest-neighbor interactions. Namely, the energy of the DH $\eta$-triplet state $\varepsilon(\eta=1,\eta^{z}=0)=-V_{\perp}$ is lower than the energy of the HH and DD pair states, $\varepsilon(\eta=1,\eta^{z}=\pm 1)=+V_{\perp}$, and is also lower than the energy of the $\eta$-singlet state $\varepsilon(\eta=0,\eta^{z}=0)=-V_{\perp}+J_{\perp}$. Therefore, it is energetically preferable that such DH pairing correlations develop as the photoexcited carriers relax into quasi-steady lowest-energy conditions. In the Supplemental Material [86], we show that DD and HH pairing correlations are indeed weaker and shorter-ranged.
Figure 2 shows the correlation functions at $n_{d}=0.2$ and $V=0.2t_{\parallel}$ in the isotropic ladder ($t_{\perp}=t_{\parallel}$). The log-log plot in Fig. 2(a) exposes that the DH pairing correlation exhibits power-law decay, while the spin, charge, and on-site pairing correlations decay exponentially, indicating that the DH pairing correlation is dominant. As shown in Fig. 2(b), the DH pairing correlation shows a sign alternation along the chain direction. The mechanism of this striped correlation will be discussed later.
To examine the pairing symmetry, following studies of chemically doped systems [57, 63, 64, 65, 66, 67], we introduce $\hat{\Delta}^{\mathrm{c}}_{j}=\sum_{\sigma,\beta=0,1}(-1)^{\beta}\left(1-\hat{n}_{j+\bar{\beta},0;\bar{\sigma}}\right)\hat{c}^{\dagger}_{j+\bar{\beta},0;\sigma}\hat{c}_{j+\beta,0;\sigma}\hat{n}_{j+\beta,0;\bar{\sigma}}/2$ for the DH pair formation along the chain direction and calculate the rung-chain DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$. As shown in Fig. 2(b), the rung-rung correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and the rung-chain correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibit opposite signs. Similar sign inversion between $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ appears in the SC correlations in chemically doped ladders, which was taken as a signature of $d$-wave pairing [57]. This intriguing finding suggests that the DH pairing correlation in the 2D square lattice would also have the $d$-wave symmetry and thus potentially bears a close similarity to the equilibrium SC state in chemically doped MIs.
In other regions of the $(n_{d},V)$ phase diagram, various phases can dominate. These include the CDW and $\eta$-pairing nonequilibrium phases, which have been previously studied in the 1D chain [43, 44]. Our results confirm their presence in the ladder geometry, see Fig. 1(a), and also reveal a spin-singlet phase distinct from the SDW phase in the 1D chain. As shown in Fig. 1(a), for $t_{\perp}=t_{\parallel}$, we observe the spin-singlet phase at $n_{d}=0$, where $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(s)}_{J}$. In this phase, the spin correlations decay exponentially [86], implying the presence of the spin gap due to the rung spin-singlet formation. Similarly to the 1D chain [43, 44], the staggered on-site pairing, i.e., $\eta$-pairing, correlations become dominant in the small $V$ region, and the CDW correlations become dominant in the large $n_{d}$ and large $V$ region [86]. The $\eta$-pairing and CDW are degenerate at $n_{d}=0.5$ and $V=0.2t_{\parallel}$. Since the system at $n_{d}=0.5$ is composed of doublons and holons without singly occupied sites, the effective model $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}$, which is equivalent to the XXZ model. The point at which the $\eta$-pairing and CDW are degenerate corresponds to the transition point between the XY and Ising phases in the XXZ model [43].
[…]
Effective model— Here, we explain the origin of the phase factor $(-1)^{r}$ in the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$. Based on the above results, we can conclude that an important aspect of DH-pairing is the formation of DH $\eta$-triplets on the spin-singlet background. Thus, we discuss these behaviors in a minimal model with local state space restricted to rung spin-singlet and rung DH $\eta$-triplet. We derive a minimal model from the local rung approximation [92], ^$\hat{\mathcal{H}}_{\rm LRA}=\left[\hat{\mathcal{H}}^{(0)}_{t_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}\right]+\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$, considering the strong rung coupling regime with the interrung hopping ^$\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$ as the perturbation. The second-order perturbation theory gives the effective Hamiltonian for the two-level system
^$\hat{\mathcal{H}}_{\rm min}=K\sum_{j}\left(\hat{\mathbf{\Delta}}_{j}\cdot\hat{\mathbf{\Delta}}_{j+1}-\frac{1}{4}\right),$ (5)
where $K=2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}+2t_{\perp})+2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}-2t_{\perp})$, see Supplemental Material for details [86]. The rung DH $\eta$-triplet and rung spin-singlet states define the effective up and down pseudospin states, respectively, see inset of Fig. 4(b). $\hat{\mathbf{\Delta}}_{j}$ is a vector composed of the pseudospin operators $\hat{\Delta}^{+}_{j}$, $\hat{\Delta}^{-}_{j}$, and $\hat{\Delta}^{z}_{j}$. As before, $\hat{\Delta}^{+}_{j}=\hat{\Delta}^{r\dagger}_{j}$ maps the rung spin-singlet state to the DH $\eta$-triplet state. The fixed number of doublons in photodoped MIs corresponds to the fixed pseudo magnetization in Eq. (5).
Figure 4: (a) Linear scale plot and (b) log-log plot of the rung-rung DH pairing correlation function with varying $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$. Here, $t_{\perp}$ is used as the unit of energy. We set $n_{d}=0.05$, $J_{\perp}=0.4t_{\perp}$, and $V_{\perp}=1.8t_{\perp}$. The interrung parameters are varied as $t_{\parallel}=(0.1+0.9p)t_{\perp}$, $J_{\parallel}=0.4pt_{\perp}$, and $V_{\parallel}=1.8pt_{\perp}$, where $0\leq p\leq 1$. Inset: correspondence between the photodoped ladder and the 1D antiferromagnetic Heisenberg model. The blue and purple ellipses represent the rung spin-singlet and rung DH $\eta$-triplet, respectively, as shown in Fig. 1(b).
This effective model assumes $V_{\perp}+J_{\perp}>|2t_{\perp}|$ to set the spin-singlet and $\eta$-triplet states to the lowest-energy configurations in the independent rung limit [86]. Under this condition, $K>0$, and therefore Eq. (5) is equivalent to the 1D antiferromagnetic Heisenberg model at a certain magnetization. The lowest-energy state of Eq. (5) coincides with the ground state of the Heisenberg model under a magnetic field. Due to the equivalence with the antiferromagnetic Heisenberg model, $\braket{\hat{\Delta}^{+}_{j+r}\hat{\Delta}^{-}_{j}}$ in Eq. (5) must have the phase factor $(-1)^{r}$, explaining the sign alternation of the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the photodoped ladder.
Going beyond the above analytical perturbative arguments, we study the stability of the sign alternation in $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{i}\hat{\Delta}^{\mathrm{r}}_{j}}$ as a function of the $t_{\parallel},J_{\parallel},V_{\parallel}$ parameters of the full model $\hat{\mathcal{H}}$ numerically. Figure 4 shows the DH pairing correlations as $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$ increase linearly, departing from ^$\hat{\mathcal{H}}_{\rm LRA}$. As expected from the effective two-level model in Eq. (5), the DH pairing correlation is well developed when different rungs are only weakly connected ($p=0$). Although the absolute value of the correlation gradually decreases at larger $p$ as the parameters approach the original $\hat{\mathcal{H}}$, the sign alternation in the correlation function remains, confirming that ^$\hat{\mathcal{H}}_{\mathrm{min}}$ captures the essential features of the pairing correlations observed in the DH pairing phase.
[…]
In contrast to the $\eta$-pairing phase that prominently appears in the $V\sim 0$ and large $n_{d}$ region, the DH pairing state can appear in the large $V$ and small $n_{d}$ region. Because strong photoexcitation to achieve high $n_{d}$ is not required and $V$ is nonzero in real materials, the DH pairing state is more experimentally accessible. For instance, ladder-type cuprates [93, 94, 95, 96] may serve as hosts for the DH pairing state. While the implementation of nearest-neighbor Coulomb interaction is a challenging ingredient for the realization with cold atoms, the first step towards this has been achieved for the Bose-Hubbard model [97].
Although photodoping is inherently a nonequilibrium phenomenon, we have approximately mapped a photodoped state to the lowest-energy state in the pseudoequilibrium condition. To further validate the emergence of the DH pairing phase and its lifetime, it would be necessary to demonstrate the time-dependent behavior of the DH pairing correlations in the optically driven Hubbard ladder using time-evolution methods [98, 99, 100, 101, 102]. These investigations remain as future challenges.
Acknowledgements.
We thank S. Ejima, Y. Murakami, S. Nishimoto, and T. Sato for fruitful discussions. This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grants No. JP20H01849, No. JP24K06939, No. JP24H00191 (T.K.), No. JP22K04907 (K.K.), and No. JP24K01333. R.U. was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program”. M.S. acknowledges the support from the UK EPSRC award under the Agreement EP/Y005090/1. Z.L. acknowledges the support by the QuantERA II JTC 2021 grant T-NiSQ by MVZI, the P1-0044 program of the Slovenian Research Agency, and ERC StG 2022 project DrumS, Grant Agreement 101077265. D.G. is supported by the Slovenian Research and Innovation Agency (ARIS) under Programs No. P1-0044, No. J1-2455, and No. MN-0016-106. Our calculations were performed using the ITensor library [103, 104].

### v2

@@
-Exotic Doublon-Holon Pairing State in Photodoped Mott Insulators
+Doublon-holon pairing state in photodoped Mott insulators
 Abstract
-We demonstrate the existence of a unique pairing state in photodoped Mott insulators on ladder geometries, characterized by quasi-long-ranged doublon-holon correlations, using the density matrix renormalization group method.
-This phase exhibits doublon-holon pairing correlations with opposite signs along the rung and chain directions, reminiscent of $d$-wave pairing in chemically doped ladder systems.
-By constructing the phase diagram, we reveal that the doublon-holon pairing state emerges between the spin-singlet phase and the charge-density-wave/$\eta$-pairing phase.
+We demonstrate the existence of an unconventional pairing state in photodoped Mott insulators on ladder and quasi-two-dimensional geometries, characterized by quasi-long-range doublon-holon correlations that signal Mott exciton condensation.
+The doublon-holon pairing exhibits correlations of $d$-wave-like symmetry, reminiscent of superconducting pairing in chemically doped Mott insulators.
+By constructing the phase diagram, using density matrix renormalization group, we reveal that the doublon-holon pairing state in the photodoped ladder emerges between the spin-singlet, charge-density-wave, and $\eta$-pairing phases.
 Our study suggests that the interplay of charge, spin, and $\eta$-spin degrees of freedom can give rise to exotic quantum many-body states in photodoped Mott insulators.
+Introduction.
 External field driving is a useful tool for generating and controlling intriguing out-of-equilibrium phenomena in correlated electron systems [1, 2, 3, 4, 5, 6].
 Photoinduced phase transitions [7, 8, 9, 10, 11, 12, 13], Floquet engineering [14, 15, 16, 17, 18], and high-harmonic generation [19, 20, 21, 22, 23, 24, 25, 26] have been extensively studied to date.
-In insulating materials, optical excitation across the band gap generates electron and hole carriers.
-When the recombination time is sufficiently long, the excited state potentially settles into a quasi-steady photodoped state, nearly conserving the number of carriers [27, 28].
-Photoexcitation in correlation-driven Mott insulators (MIs) creates two types of carriers, doublons (doubly occupied sites) and holons (empty sites), see Fig.
-1(b) [29, 30, 6, 31, 32, 33, 34], resulting in a unique photodoped state.
-In a large-gap MI, the carriers are expected to have long lifetimes, leading to a quasi-steady state where doublons and holons coexist [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
-Figure 1: (a) Phase diagram of the photodoped ladder-type MI as a function of the doublon density $n_{d}$ and the intersite Coulomb interaction $V$ with the blue, red, green, and purple points representing the predominance of the spin, $\eta$-pairing, CDW, and doublon-holon (DH) pairing correlations, respectively.
+In semiconductors, optical excitation across the band gap generates electron and hole carriers.
+Similarly, photoexcitation in correlation-driven Mott insulators (MIs) creates two types of carriers, doublons (doubly occupied sites) and holons (empty sites) [see Fig.
+1(b)] [27, 28, 6, 29, 30, 31], resulting in a photodoped state distinct from conventional photoexcited semiconductors.
+In a large-gap MI, the carriers are expected to have long lifetimes, leading to a quasisteady state where doublons and holons coexist [32, 33, 34, 35, 36, 37, 38, 39, 40, 41] and possibly bind in excitons [42, 43, 36, 37, 38, 44], which are distinct from excitons in photodoped semiconductors [45, 46].
+Figure 1: (a) Phase diagram of the photodoped ladder as a function of the doublon density $n_{d}$ and the intersite Coulomb interaction $V$ with the blue, red, green, and purple points representing the predominance of the spin, $\eta$-pairing, CDW, and doublon-holon (DH) pairing correlations, respectively.
 At the orange point ($n_{d}=0.5$ and $V=0.2t_{\parallel}$), the $\eta$-pairing and CDW states are degenerate.
@@
 (b) DH pairing state in the photodoped ladder with doublons and holons generated by light.
-The DH pairing correlation is characterized by an operator that maps spin singlets on nearest-neighbor sites to DH $\eta$-triplets.
-Photodoping can be a trigger for creating exotic many-body states that are inaccessible in equilibrium, such as the $\eta$-pairing state [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
-The photodoped states have been investigated in detail in the one-dimensional (1D) Hubbard chain, where the phase diagram comprises the spin density wave (SDW) phase at low doublon density $n_{d}$ and the charge density wave (CDW) and $\eta$-pairing phases from intermediate to high $n_{d}$ [43, 44].
+(c) Spin interaction that acts between up and down spins, and $\eta$-spin interaction that acts between a doublon and a holon.
+(d) Operator that maps a spin singlet on nearest-neighbor sites to a DH $\eta$ triplet.
+Photodoping can be a trigger for creating exotic many-body states that are inaccessible in equilibrium, such as the $\eta$-pairing state, which is characterized by long-range correlation of on-site electronic pairs [47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
+The photodoped states have been theoretically investigated in detail in the one-dimensional (1D) Hubbard chain [40, 41] and in high-dimensional Hubbard systems [57, 58, 59].
+In the former case, the phase diagram comprises the spin-density-wave (SDW) phase at low doublon density $n_{d}$ and the charge-density-wave (CDW) and $\eta$-pairing phases from intermediate to high $n_{d}$ [40, 41].
 Although even richer physical properties are anticipated in the two-dimensional (2D) systems because of the increased geometric degrees of freedom and the lack of spin-charge-$\eta$ separation [41], these properties of photodoped 2D MIs are yet to be understood.
 Ladder geometry can serve as a bridge between 1D and 2D systems and is more manageable for numerical approaches.
-For example, density matrix renormalization group (DMRG) calculations, which enable us to obtain numerically precise low-dimensional wave functions, have shown that chemically hole-doped ladders exhibit pair correlations corresponding to $d$-wave superconductivity (SC) in a 2D square lattice [57, 58, 59, 60, 61].
+For example, density matrix renormalization group (DMRG) calculations, which enable us to obtain numerically precise low-dimensional wave functions, have shown that chemically hole-doped ladders exhibit pair correlations corresponding to $d$-wave superconductivity (SC) in a 2D square lattice [60, 61, 62, 63, 64, 65].
 These studies are an important stepping stone to research on $d$-wave SC in the Hubbard and $t$-$J$ models on multileg cylinders, mimicking the 2D square lattice of cuprate superconductors [66, 67, 68, 69, 70, 71].
 Furthermore, the pairing mechanism in ladder and bilayer systems [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] has attracted significant attention associated with the recent discovery of high-temperature SC in bilayer nickelates [84, 85, 86, 87].
-Because a strong interchain spin coupling favors the formation of spin-singlet rungs, hole doping to the ladder system leads to the formation of interchain hole-hole pairs to minimize the disruption of spin-singlet bonds, resulting in the development of SC correlations [68].
-Realization of such a pairing mechanism has been recently proposed also with cold atoms in optical lattices [84].
-From these considerations, we arrive at the crucial question: what are the equivalent paired states in photodoped MIs?
-In this Letter, we show that photodoped MIs in the ladder geometry and quasi-stationary situation exhibit an exotic pairing state composed of doublons and holons (Fig.
+Because a strong interchain spin coupling favors the formation of spin-singlet rungs, hole doping to the ladder system leads to the formation of interchain hole-hole pairs to minimize the disruption of spin-singlet bonds [72].
+Realization of such a pairing mechanism has also been recently proposed with cold atoms in optical lattices [88].
+From these considerations, we arrive at the crucial question: what are the equivalent correlation-driven paired states in photodoped MIs?
+In this Letter, we show that photodoped MIs in a quasistationary situation exhibit a pairing state composed of doublons and holons (Fig.
 1).
-This pairing is characterized by the quasi-long-range correlation of doublon-holon (DH) pairs on the spin background shown in Fig.
-1(b).
-We find opposite phases for the rung and chain pairing, analogous to signatures of $d$-wave pairing in chemically doped ladders.
-We reveal the roles of interchain coupling and propose an effective minimal model that can capture the pairing properties.
-Model and Methods—To model MIs in the presence of nonlocal interactions, we start with the extended Hubbard Hamiltonian in a ladder geometry
-^Hub=subscript^$\hat{\mathcal{H}}_{\mathrm{Hub}}=$ $-t_{\parallel}\sum_{j,\alpha,\sigma}\left(\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j+1,\alpha;\sigma}+\mathrm{H.c.}\right)$
-$-t_{\perp}\sum_{j,\sigma}\left(\hat{c}^{\dagger}_{j,0;\sigma}\hat{c}_{j,1;\sigma}+\mathrm{H.c.}\right)+U\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha;\uparrow}\hat{n}^{\prime}_{j,\alpha;\downarrow}$
-$+V_{\parallel}\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha}\hat{n}^{\prime}_{j+1,\alpha}+V_{\perp}\sum_{j}\hat{n}^{\prime}_{j,0}\hat{n}^{\prime}_{j,1}.$ (1)
-Here, $\hat{c}^{\dagger}_{j,\alpha;\sigma}$ ($\hat{c}_{j,\alpha;\sigma}$) is the creation (annihilation) operator for a fermion with spin $\sigma=\uparrow$, $\downarrow$ at site $j$ on chain $\alpha=0,1$.
-$\hat{n}_{j,\alpha;\sigma}=\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j,\alpha;\sigma}$, $\hat{n}^{\prime}_{j,\alpha;\sigma}=\hat{n}_{j,\alpha;\sigma}-1/2$, and $\hat{n}^{\prime}_{j,\alpha}=\hat{n}^{\prime}_{j,\alpha;\uparrow}+\hat{n}^{\prime}_{j,\alpha;\downarrow}$.
-$t_{\parallel}$ and $t_{\perp}$ are the hopping integrals along the chain and rung directions, respectively.
-$U>0$ is the on-site Coulomb interaction.
-$V_{\parallel}>0$ and $V_{\perp}>0$ are the nearest-neighbor Coulomb interactions along the chain and rung directions, respectively.
+In the ladder geometry, this pairing state is characterized by the quasi-long-range correlation of doublon-holon (DH) pairs (Mott excitons), suggesting exciton condensation.
+Due to strong correlations, the state is distinct from excitonic condensates in semiconductors.
+Namely, we observe opposite phases for the rung and chain pairing, analogous to signatures of superconducting $d$-wave pairing in chemically doped ladders.
+These findings are further corroborated by results on a multileg cylinder mimicking the 2D square lattice.
+Model. To model MIs in the presence of nearest-neighbor interactions, we start with the standard extended Hubbard Hamiltonian
+$\hat{\mathcal{H}}_{\mathrm{Hub}}=\sum_{\langle\bm{ij}\rangle}\hat{T}_{\bm{ij}}+U\sum_{\bm{j}}\hat{n}^{\prime}_{\bm{j};\uparrow}\hat{n}^{\prime}_{\bm{j};\downarrow}+\sum_{\langle\bm{ij}\rangle}V_{\bm{ij}}\hat{n}^{\prime}_{\bm{i}}\hat{n}^{\prime}_{\bm{j}},$ (1)
+where $\hat{T}_{\bm{ij}}=-t_{\bm{ij}}\sum_{\sigma}\hat{c}^{\dagger}_{\bm{i};\sigma}\hat{c}_{\bm{j};\sigma}+{\rm H.c.}$ is the hopping term.
+$\hat{c}^{\dagger}_{\bm{j};\sigma}$ ($\hat{c}_{\bm{j};\sigma}$) is the fermion creation (annihilation) operator at site $\bm{j}$ for spin $\sigma=\uparrow,$ $\downarrow$.
+$\hat{n}_{\bm{j};\sigma}=\hat{c}^{\dagger}_{\bm{j};\sigma}\hat{c}_{\bm{j};\sigma}$, $\hat{n}^{\prime}_{\bm{j};\sigma}=\hat{n}_{\bm{j};\sigma}-1/2$, and $\hat{n}^{\prime}_{\bm{j}}=\hat{n}^{\prime}_{\bm{j};\uparrow}+\hat{n}^{\prime}_{\bm{j};\downarrow}$.
+Summation $\braket{\bm{i}\bm{j}}$ runs over nearest-neighbor sites.
+In the two-leg ladder, $\bm{j}=(j,\alpha)$ denotes site at position $j$ on chain $\alpha$ [=0=0 (lower), $1$ (upper)], and we distinguish the hopping integrals $t_{\bm{ij}}=t_{\parallel},t_{\perp}$ and the nearest-neighbor Coulomb interaction $V_{\bm{ij}}=V_{\parallel},V_{\perp}$ along the chain and rung directions, respectively.
+In the 2D square lattice, $t_{\bm{ij}}=t_{\rm h}$ and $V_{\bm{ij}}=V$.
+$U$ is the on-site Coulomb repulsion.
+The lattice constant $a$ is set to one, and we use $t_{\parallel}$ in the two-leg ladder ($t_{\rm h}$ in the square lattice) as the unit of energy unless otherwise specified.
 In photodoped MIs, doublons and holons are generated by an optical excitation, as shown in Fig.
 1(b).
-When $U\gg t_{\parallel},t_{\perp}$, the recombination time is sufficiently long [35, 36, 37, 38, 39, 41, 40, 42] and the system can be considered to be in a pseudoequilibrium state where doublons and holons are conserved [43, 44].
-Assuming $U\gg t_{\parallel},t_{\perp}$, we introduce an effective model derived by the Schrieffer-Wolff transformation that excludes the hopping terms changing the number of doublons and holons [39, 41, 43, 85].
-The Hamiltonian of the effective $t$-$J$-$V$ model is given by
-^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}^{(0)}_{t}+\hat{\mathcal{H}}^{(s)}_{J}+\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}.$ (2)
-Here, ^$\hat{\mathcal{H}}^{(0)}_{t}=\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}+\hat{\mathcal{H}}^{(0)}_{t_{\perp}}$ denotes the doublon-number-conserving hopping, see Supplemental Material for details [86], and ^$\hat{\mathcal{H}}_{V}=\hat{\mathcal{H}}_{V_{\parallel}}+\hat{\mathcal{H}}_{V_{\perp}}$ represents the nearest-neighbor Coulomb interactions $V_{\parallel}$ and $V_{\perp}$ in Eq.
-(1).
-^$\hat{\mathcal{H}}^{(s)}_{J}=\hat{\mathcal{H}}^{(s)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}$ describes the Heisenberg-type spin-exchange interaction, while ^$\hat{\mathcal{H}}^{(\eta)}_{J}=\hat{\mathcal{H}}^{(\eta)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}$ represents the $\eta$-spin interaction for doublons and holons [43, 85].
-These nearest-neighbor interactions between sites $\bm{i}$ and $\bm{j}$ are given by
-^$\hat{\mathcal{H}}^{(s)}_{J_{\bm{i}\bm{j}}}=J_{\bm{i}\bm{j}}\left(\hat{\bm{s}}_{\bm{i}}\cdot\hat{\bm{s}}_{\bm{j}}-\frac{1}{4}\delta_{1,\hat{n}_{\bm{i}}\hat{n}_{\bm{j}}}\right),$ (3)
-^$\hat{\mathcal{H}}^{(\eta)}_{J_{\bm{i}\bm{j}}}=-J_{\bm{i}\bm{j}}\left(\hat{\bm{\eta}}_{\bm{i}}\cdot\hat{\bm{\eta}}_{\bm{j}}-\frac{1}{4}\left(1-\delta_{1,\hat{n}_{\bm{i}}}\right)\left(1-\delta_{1,\hat{n}_{\bm{j}}}\right)\right),$ (4)
-where $\bm{j}$ denotes $(j,\alpha)$.
-The spin operator is defined as $\hat{\bm{s}}_{\bm{j}}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{j,\alpha;\sigma}\bm{\sigma}_{\sigma,\sigma^{\prime}}\hat{c}_{j,\alpha;\sigma^{\prime}}/2$, where $\bm{\sigma}$ is the vector of Pauli matrices.
-The $\eta$-spin operator $\hat{\bm{\eta}}_{\bm{j}}$ is given by $\hat{\eta}^{+}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}^{\dagger}_{j,\alpha;\downarrow}\hat{c}^{\dagger}_{j,\alpha;\uparrow}$, $\hat{\eta}^{-}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$, and $\hat{\eta}^{z}_{\bm{j}}=\left(\hat{n}_{j,\alpha}-1\right)/2$ [87].
-The explicit Hamiltonian for the ladder system is presented in the Supplemental Material [86].
-For the hopping integral $t_{\bm{i}\bm{j}}$, the coupling constant is characterized by $J_{\bm{i}\bm{j}}=4t^{2}_{\bm{i}\bm{j}}/U>0$ (at $V_{\bm{i}\bm{j}}=0$).
-The $V$ term takes the form $V_{\bm{i}\bm{j}}\hat{n}^{\prime}_{\bm{i}}\hat{n}^{\prime}_{\bm{j}}=4V_{\bm{i}\bm{j}}\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}$, thereby introducing Ising-type anisotropy of $\eta$-spin.
-In contrast to the chemically doped $t$$J$ model, $\hat{\mathcal{H}}$ includes the interaction between doublons and holons described by $\eta$-spin, which enables the formation of $\eta$-pairing states in photodoped MIs [51, 43, 44].
-We describe the photodoped ladder as the lowest-energy state of $\hat{\mathcal{H}}$ with a fixed doublon number.
-We obtain the state using the DMRG method [88, 89, 90].
-Unless otherwise specified, we use $t_{\parallel}$ as the unit of energy and set the chain length to $L=160$ with open boundary conditions.
-The total number of sites is $2L$, and the numbers of up- and down-spin fermions are set to $N_{\uparrow}=N_{\downarrow}=L$ (i.e., half-filling without spin polarization), while the state is configured as a function of the doublon number.
+Their recombination rate $\Gamma_{\rm R}\propto\exp(-U/\epsilon_{0})$ is exponentially suppressed with the number of excitations that are emitted in the recombination process [32, 33, 34, 35, 36, 37, 39, 38].
+For low densities of doublons and holons, the recombination predominantly happens via emission of magnetic processes, $\epsilon_{0}$ being their characteristic energy [44].
+For higher densities, scattering with other doublons and holons ($\epsilon_{0}\sim t_{\bm{ij}}$) can also contribute [33, 34].
+Thus, when $U\gg\epsilon_{0}$, the system can be on the timescales that are parametrically long in $U$ considered to be in a pseudoequilibrium state where doublons and holons are conserved [40, 41].
+To efficiently consider the case $U\gg\epsilon_{0}$, we introduce an effective model derived by the Schrieffer-Wolff transformation that conserves the number of doublons and holons [36, 37, 40, 89] with the dominant (lowest order) terms:
+$\hat{\mathcal{H}}$= $-\sum_{\langle\bm{ij}\rangle,\sigma}t_{\bm{ij}}\left(\hat{d}^{\dagger}_{\bm{i};\sigma}\hat{d}_{\bm{j};\sigma}-\hat{h}^{\dagger}_{\bm{i};\sigma}\hat{h}_{\bm{j};\sigma}+{\rm H.c.}\right)$
+$+\sum_{\langle\bm{ij}\rangle}J_{\bm{i}\bm{j}}\left(\hat{\bm{s}}_{\bm{i}}\cdot\hat{\bm{s}}_{\bm{j}}-\frac{1}{4}\delta_{1,\hat{n}_{\bm{i}}}\delta_{1,\hat{n}_{\bm{j}}}\right)$ (2)
+$-\sum_{\langle\bm{ij}\rangle}\left[J_{\bm{i}\bm{j}}\left(\hat{\bm{\eta}}_{\bm{i}}\cdot\hat{\bm{\eta}}_{\bm{j}}-\frac{1}{4}\bar{\delta}_{1,\hat{n}_{\bm{i}}}\bar{\delta}_{1,\hat{n}_{\bm{j}}}\right)-4V_{\bm{i}\bm{j}}\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}\right].$
+The first term denotes the hopping of doublons and holons, written explicitly with doublon $\hat{d}^{\dagger}_{\bm{j};\sigma}=\hat{c}^{\dagger}_{\bm{j};\sigma}\hat{n}_{\bm{j};\bar{\sigma}}$ and holon $\hat{h}^{\dagger}_{\bm{j};\sigma}=\hat{c}_{\bm{j};\sigma}(1-\hat{n}_{\bm{j};\bar{\sigma}})$ creation operators [36], where $\bar{\sigma}=\downarrow$ $(\uparrow)$ for $\sigma=\uparrow$ $(\downarrow)$.
+The second term describes the SU(2)(2) symmetric spin-exchange interaction, which is also considered in the traditional $t$-$J$ model.
+The spin operator is defined as $\hat{\bm{s}}_{\bm{j}}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{\bm{j};\sigma}\bm{\sigma}_{\sigma\sigma^{\prime}}\hat{c}_{\bm{j};\sigma^{\prime}}/2$, where $\bm{\sigma}$ is the vector of Pauli matrices.
+For the hopping $t_{\bm{i}\bm{j}}$, the coupling constant is given by $J_{\bm{i}\bm{j}}=4t^{2}_{\bm{i}\bm{j}}/U>0$ (at $V_{\bm{i}\bm{j}}=0$).
+Similarly, the third term represents the interaction between doublons and holons [40, 89].
+In the absence of $V_{\bm{i}\bm{j}}$, this term reveals the SU(2)(2) symmetry of the Hubbard model, represented by the $\eta$ spin.
+The $\eta$-spin operator $\hat{\bm{\eta}}_{\bm{j}}$ is given by $\hat{\eta}^{+}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}^{\dagger}_{\bm{j};\downarrow}\hat{c}^{\dagger}_{\bm{j};\uparrow}$, $\hat{\eta}^{-}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}_{\bm{j};\uparrow}\hat{c}_{\bm{j};\downarrow}$, and $\hat{\eta}^{z}_{\bm{j}}=\left(\hat{n}_{\bm{j}}-1\right)/2$ [90].
+In Eq.
+(Doublon-holon pairing state in photodoped Mott insulators), $\bar{\delta}_{1,\hat{n}_{\bm{j}}}=1-\delta_{1,\hat{n}_{\bm{j}}}$, and the $\eta$ term acts only on empty ($\circ$) and doubly occupied ($\uparrow\downarrow$) sites.
+$\hat{\eta}^{+}_{\bm{i}}\hat{\eta}^{-}_{\bm{j}}$ swaps the positions of neighboring doublon and holon, while $\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}$ measures the energy according to the occupancy status of sites $\bm{i}$ and $\bm{j}$ [91].
+The $\eta$-spin interaction in Eq.
+(Doublon-holon pairing state in photodoped Mott insulators) is ferromagnetic, which preferably forms an $\eta$-spin triplet state at two sites.
+This $\eta$-spin interaction is activated by photodoping, while it is negligible in the traditional $t$-$J$ model for chemically doped systems.
+As is clear from Eq.
+(Doublon-holon pairing state in photodoped Mott insulators), $V_{\bm{i}\bm{j}}$ manifestly breaks the $\eta$-SU(2)(2) symmetry and leads to an anisotropic XXZ-type Hamiltonian for the $\eta$ sector.
+Overall, $\hat{\mathcal{H}}$ is expected to describe metastable phases in photodoped MIs with a similar level of rigor as the well-known $t$-$J$ model can describe chemically doped materials.
+For $U\gg t_{\bm{ij}},J_{\bm{ij}}$, implying slow recombination of doublons and holons, photoexcited carriers have time to relax to the bottom of the upper band.
+Therefore, we can transiently describe the photodoped state as the lowest-energy state of $\hat{\mathcal{H}}$ at a fixed doublon density $n_{d}$ that corresponds to the excitation protocol.
+In the two-leg ladder of length $L$, the doublon density is given by $n_{d}=\sum_{j,\alpha}\braket{\hat{n}_{j,\alpha;\uparrow}\hat{n}_{j,\alpha;\downarrow}}/(2L)$.
+The lowest-energy state, representing the transient photodoped state, is obtained with the DMRG method [92, 93, 94] within a sector with a fixed number of doublons [95].
+In the two-leg ladder, the total number of sites is $N_{\rm s}=2L$.
+The numbers of up- and down-spin fermions are set to $N_{\uparrow}=N_{\downarrow}=N_{\rm s}/2$ (i.e., half filling without spin polarization), while the state is configured as a function of the doublon number.
+The given doublon number $N_{d}$ determines the doublon density as $n_{d}=N_{d}/N_{\rm s}$.
+We employ open boundary conditions along the chain ($x$) direction.
+In the two-leg ladder, we set the chain length to $L=160$ unless otherwise specified.
 The bond dimension is set to $m=10000$ with the maximum truncation error typically being on the order of $1.0\times 10^{-7}$.
-Numerical Results—To investigate which nonequilibrium phases can emerge in photodoped ladders, we study the behavior of correlation functions $C(r)=\braket{\hat{O}^{\dagger}_{j_{0}+r}\hat{O}_{j_{0}}}$ between operators $\hat{O}_{j}$ that are $r$ sites apart.
-We calculate correlations with respect to the site $j_{0}=L/4+1$ to minimize open boundary effects.
-The dominant nonequilibrium phase is determined by identifying the correlation function that exhibits the slowest decay with spatial distance $r$.
-Because the correlations between the single-site operators are symmetric on both chains, we consider the $\alpha=0$ chain index in the operators.
-We use $\hat{O}_{j}=\hat{n}_{j,\alpha;\uparrow}-\hat{n}_{j,\alpha;\downarrow}$ for the spin correlation, $\hat{O}_{j}=\hat{n}_{j,\alpha}-1$ for the charge correlation, and $\hat{O}_{j}=\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$ for the on-site pairing correlation.
+For the 2D square lattice, we employ an $8\times 4$ site cylinder with periodic boundary conditions along the $y$ direction, where the bond dimension and the maximum truncation error are $m=15000$ and $1.0\times 10^{-5}$, respectively.
+To identify nonequilibrium phases in photodoped states, we analyze correlation functions $C(r)=\braket{\hat{O}^{\dagger}_{j_{0}+r}\hat{O}_{j_{0}}}$ between operators $\hat{O}_{j}$ separated by $r$ sites.
+The reference site is fixed at $j_{0}=L/4+1$ with open boundary conditions.
+We present the site-averaged correlation functions in the Supplemental Material [91] and confirm that our conclusion does not rely on the definition of the correlation function.
+The dominant phase is determined by the correlation function that exhibits the slowest decay with distance $r$.
 Figure 2: (a) Log-log plot and (b) linear scale plot of the correlation functions for $n_{d}=0.2$, $t_{\perp}=t_{\parallel}$, $J_{\perp}=J_{\parallel}=0.4t_{\parallel}$, and $V=V_{\perp}=V_{\parallel}=0.2t_{\parallel}$.
@@
 The on-site pairing correlations are multiplied by two for comparison on the same scale.
-The central result of this Letter is the finding of the DH pairing phase in a finite region of doublon density $n_{d}=\sum_{j,\alpha}\braket{\hat{n}_{j,\alpha;\uparrow}\hat{n}_{j,\alpha;\downarrow}}/(2L)$ and at sufficiently strong nonlocal repulsion $V$, shown in Fig.
-1(a).
-This DH pairing state is characterized by the correlation function $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ with $\hat{\Delta}^{\mathrm{r}}_{j}=\sum_{\alpha,\sigma}(-1)^{\alpha}\left(1-\hat{n}_{j,\bar{\alpha};\bar{\sigma}}\right)\hat{c}^{\dagger}_{j,\bar{\alpha};\sigma}\hat{c}_{j,\alpha;\sigma}\hat{n}_{j,\alpha;\bar{\sigma}}/2$, where $\bar{\alpha}=1$ $(0)$ for $\alpha=0$ $(1)$, and $\bar{\sigma}=\downarrow$ ()($\uparrow$)( ↑ ) for $\sigma=\uparrow$ ()($\downarrow$)( ↓ ).
-The operator $\hat{\Delta}^{\mathrm{r}\dagger}_{j}$ creates a rung $\eta$-triplet with a doublon and a holon on the rung spin-singlet background, as shown in Fig.
-1(b) [91].
-Here, the $\eta$-triplet states ($\eta=1$) with $\eta^{z}=-1,0$, and $+1$ correspond to holon-holon (HH), DH, and doublon-doublon (DD) pair states, respectively [86].
-Considering ^$\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}$ on a single rung, we can see that the formation of the rung DH pair in the $\eta$-triplet state is energetically favorable in the presence of the nearest-neighbor interactions.
-Namely, the energy of the DH $\eta$-triplet state $\varepsilon(\eta=1,\eta^{z}=0)=-V_{\perp}$ is lower than the energy of the HH and DD pair states, $\varepsilon(\eta=1,\eta^{z}=\pm 1)=+V_{\perp}$, and is also lower than the energy of the $\eta$-singlet state $\varepsilon(\eta=0,\eta^{z}=0)=-V_{\perp}+J_{\perp}$.
-Therefore, it is energetically preferable that such DH pairing correlations develop as the photoexcited carriers relax into quasi-steady lowest-energy conditions.
-In the Supplemental Material [86], we show that DD and HH pairing correlations are indeed weaker and shorter-ranged.
-Figure 2 shows the correlation functions at $n_{d}=0.2$ and $V=0.2t_{\parallel}$ in the isotropic ladder ($t_{\perp}=t_{\parallel}$).
+Numerical results.
+The central result of this Letter is the finding of a DH pairing phase, characterized with the slowest power-law decay in the correlation functions $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$, defined via the operators
+$\hat{\Delta}^{\mathrm{\mathrm{r}\dagger}}_{j}$ $=\frac{1}{2}\sum_{\alpha=0,1}\sum_{\sigma}(-1)^{\alpha}\hat{d}^{\dagger}_{j,\alpha;\sigma}\hat{h}^{\dagger}_{j,\bar{\alpha};\sigma},$ (3)
+$\hat{\Delta}^{\mathrm{\mathrm{c}\dagger}}_{j}$ $=\frac{1}{2}\sum_{\beta=0,1}\sum_{\sigma}(-1)^{\beta}\hat{d}^{\dagger}_{j+\beta,0;\sigma}\hat{h}^{\dagger}_{j+\bar{\beta},0;\sigma},$ (4)
+where $\bar{\alpha}=1$ $(0)$ for $\alpha=0$ $(1)$.
+These operators act on nearest-neighboring sites that are in the spin-singlet state and transfer it to the state with a doublon and a holon in the $\eta$-triplet state, as shown in Fig.
+1(d), along the rung (r) and chain (c) directions, respectively.
+The definitions of the pair operators given above correspond to the formation of a local component of a Mott exciton, energetically favored by the interactions between nearest-neighbor sites (see Supplemental Material [91] for details).
+Figure 2 shows the correlation functions at $n_{d}=0.2$ and $V=0.2t_{\parallel}$ in the isotropic ladder ($t_{\perp}=t_{\parallel}$, $V_{\perp}=V_{\parallel}=V$).
 The log-log plot in Fig.
-2(a) exposes that the DH pairing correlation exhibits power-law decay, while the spin, charge, and on-site pairing correlations decay exponentially, indicating that the DH pairing correlation is dominant.
-As shown in Fig.
-2(b), the DH pairing correlation shows a sign alternation along the chain direction.
-The mechanism of this striped correlation will be discussed later.
-To examine the pairing symmetry, following studies of chemically doped systems [57, 63, 64, 65, 66, 67], we introduce $\hat{\Delta}^{\mathrm{c}}_{j}=\sum_{\sigma,\beta=0,1}(-1)^{\beta}\left(1-\hat{n}_{j+\bar{\beta},0;\bar{\sigma}}\right)\hat{c}^{\dagger}_{j+\bar{\beta},0;\sigma}\hat{c}_{j+\beta,0;\sigma}\hat{n}_{j+\beta,0;\bar{\sigma}}/2$ for the DH pair formation along the chain direction and calculate the rung-chain DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$.
+2(a) shows that the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibits a power-law decay, while the spin ($\hat{O}_{j}=\hat{n}_{j,\alpha;\uparrow}-\hat{n}_{j,\alpha;\downarrow}$), charge ($\hat{O}_{j}=\hat{n}_{j,\alpha}-1$), and on-site pair ($\hat{O}_{j}=\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$) correlations [96] decay exponentially, indicating that the DH pairing correlation is dominant.
 As shown in Fig.
 2(b), the rung-rung correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and the rung-chain correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibit opposite signs.
-Similar sign inversion between $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ appears in the SC correlations in chemically doped ladders, which was taken as a signature of $d$-wave pairing [57].
-This intriguing finding suggests that the DH pairing correlation in the 2D square lattice would also have the $d$-wave symmetry and thus potentially bears a close similarity to the equilibrium SC state in chemically doped MIs.
+Similar sign inversion between superconducting rung-rung and rung-chain correlators appears in chemically doped ladders [91], which was taken as a signature of $d$-wave pairing [61].
+This intriguing finding suggests that the DH pairing correlation in 2D systems potentially bears a close similarity to the equilibrium $d$-wave SC state in chemically doped MIs.
+[…]
+The emergence of long-range correlations along the chain is a nontrivial result that cannot be explained by local pair formation alone.
+Our DMRG calculations demonstrate that the DH pairing correlations can grow over long distances, revealing not only local exciton formation but also a long-range development of the pairing correlation.
+The DH pairing state is stable in a wide regime of doublon density $n_{d}$ and Coulomb interaction $V$ as shown in Fig.
+1(a) [97].
 In other regions of the $(n_{d},V)$ phase diagram, various phases can dominate.
-These include the CDW and $\eta$-pairing nonequilibrium phases, which have been previously studied in the 1D chain [43, 44].
-Our results confirm their presence in the ladder geometry, see Fig.
-1(a), and also reveal a spin-singlet phase distinct from the SDW phase in the 1D chain.
-As shown in Fig.
-1(a), for $t_{\perp}=t_{\parallel}$, we observe the spin-singlet phase at $n_{d}=0$, where $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(s)}_{J}$.
-In this phase, the spin correlations decay exponentially [86], implying the presence of the spin gap due to the rung spin-singlet formation.
-Similarly to the 1D chain [43, 44], the staggered on-site pairing, i.e., $\eta$-pairing, correlations become dominant in the small $V$ region, and the CDW correlations become dominant in the large $n_{d}$ and large $V$ region [86].
-The $\eta$-pairing and CDW are degenerate at $n_{d}=0.5$ and $V=0.2t_{\parallel}$.
-Since the system at $n_{d}=0.5$ is composed of doublons and holons without singly occupied sites, the effective model $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}$, which is equivalent to the XXZ model.
-The point at which the $\eta$-pairing and CDW are degenerate corresponds to the transition point between the XY and Ising phases in the XXZ model [43].
-[…]
-[…]
-Effective model— Here, we explain the origin of the phase factor $(-1)^{r}$ in the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$.
-Based on the above results, we can conclude that an important aspect of DH-pairing is the formation of DH $\eta$-triplets on the spin-singlet background.
-Thus, we discuss these behaviors in a minimal model with local state space restricted to rung spin-singlet and rung DH $\eta$-triplet.
-We derive a minimal model from the local rung approximation [92], ^$\hat{\mathcal{H}}_{\rm LRA}=\left[\hat{\mathcal{H}}^{(0)}_{t_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}\right]+\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$, considering the strong rung coupling regime with the interrung hopping ^$\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$ as the perturbation.
+Their detailed properties are presented in the Supplemental Material [91].
+Similar to the 1D chain [40, 41], the staggered on-site pairing, i.e., $\eta$-pairing, correlations become dominant in the small $V$ region, and the CDW correlations become dominant in the large $n_{d}$ and large $V$ region.
+At $n_{d}=0$, we observe the spin-singlet phase, which is distinct from the SDW phase in the 1D chain, because a coupled spin chain with a spin gap [98] is stabilized.
+Figure 3: (a) Rung-rung DH pairing correlation function for various $t_{\perp}/t_{\parallel}$ values at $L=120$, $n_{d}=0.1$ and $V=V_{\perp}=V_{\parallel}=0.4t_{\parallel}$, where $J_{\perp}/J_{\parallel}=(t_{\perp}/t_{\parallel})^{2}$ with $J_{\parallel}=0.4t_{\parallel}$.
+(b) Linear scale and (c) log-log plots of the DH pairing correlation functions in the four-leg cylinder, where $n_{d}=0.125$, $J=0.4t_{\rm h}$, and $V=0.8t_{\rm h}$.
+To examine the evolution of pairing tendency from 1D to ladder geometry, Fig.
+3(a) shows the DH pairing correlations for various $t_{\perp}/t_{\parallel}$ values.
+The DH pairing correlations increase as $t_{\perp}/t_{\parallel}$ increases.
+While the SDW correlation is dominant in the small-$t_{\perp}$ regime [91], the DH pairing correlation becomes dominant at $t_{\perp}/t_{\parallel}\gtrsim 1$.
+This $t_{\perp}$ dependence suggests that DH pairing does not arise from one-dimensionality.
+To extrapolate our observation towards 2D systems, we compute the DH pairing correlation on a cylinder square lattice with $8\times 4$ sites and periodic boundary conditions in the $y$ direction.
+Figures 3(b) and 3(c) present the DH pairing correlations calculated by the DMRG method.
+Results on a cylinder are consistent with those on a ladder: we find a slow power-law decay of the pairing correlation, accompanied by the sign alternations.
+This suggests that DH pairing should not be limited to the ladder system and should persist in 2D.
+Effective model.
+Having observed a stable staggered nature of the DH pairing, we derive here a minimal model that can explain the phase factor of $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the ladder setup.
+The $t_{\perp}/t_{\parallel}$ dependence shown in Fig.
+3(a) suggests that a background of rung spin singlets, stabilized by the strong interchain coupling via $t_{\perp}$, is favorable for the emergence of the DH pairing phase on top of it.
+The preferable formation of rung spin singlets as $t_{\perp}/t_{\parallel}$ increases is confirmed in the Supplemental Material [91].
+Since the formation of DH $\eta$ triplets on the spin-singlet background is important for DH pairing characterized by the $\hat{\Delta}^{r}_{j}$ operator [see Fig.
+1(d)], we discuss these behaviors in an effective model with local state space restricted to rung spin singlet and rung DH $\eta$ triplet.
+We derive a minimal model from the local rung approximation [99], $\hat{\mathcal{H}}_{\rm LRA}=\hat{\mathcal{H}}_{\perp}+\hat{\mathcal{T}}_{\parallel}$, considering the strong rung coupling in $\hat{\mathcal{H}}_{\perp}$ [rung-direction component in Eq.
+(Doublon-holon pairing state in photodoped Mott insulators)] with the interrung hopping $\hat{\mathcal{T}}_{\parallel}$ as the perturbation.
 The second-order perturbation theory gives the effective Hamiltonian for the two-level system
-^$\hat{\mathcal{H}}_{\rm min}=K\sum_{j}\left(\hat{\mathbf{\Delta}}_{j}\cdot\hat{\mathbf{\Delta}}_{j+1}-\frac{1}{4}\right),$ (5)
-where $K=2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}+2t_{\perp})+2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}-2t_{\perp})$, see Supplemental Material for details [86].
-The rung DH $\eta$-triplet and rung spin-singlet states define the effective up and down pseudospin states, respectively, see inset of Fig.
-4(b).
+$\hat{\mathcal{H}}_{\rm min}=K\sum_{j}\left(\hat{\mathbf{\Delta}}_{j}\cdot\hat{\mathbf{\Delta}}_{j+1}-\frac{1}{4}\right),$ (5)
+where $K=2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}+2t_{\perp})+2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}-2t_{\perp})$ (see Supplemental Material [91] for details).
+The rung DH $\eta$-triplet and rung spin-singlet states define the effective up and down pseudospin states, respectively [see the inset of Fig.
+4(b)].
 $\hat{\mathbf{\Delta}}_{j}$ is a vector composed of the pseudospin operators $\hat{\Delta}^{+}_{j}$, $\hat{\Delta}^{-}_{j}$, and $\hat{\Delta}^{z}_{j}$.
 As before, $\hat{\Delta}^{+}_{j}=\hat{\Delta}^{r\dagger}_{j}$ maps the rung spin-singlet state to the DH $\eta$-triplet state.
-The fixed number of doublons in photodoped MIs corresponds to the fixed pseudo magnetization in Eq.
+The fixed number of doublons in a photodoped MI corresponds to the fixed pseudomagnetization in Eq.
 (5).
+Since this effective model is derived under the condition $K>0$ [91], Eq.
+(5) is equivalent to the 1D antiferromagnetic Heisenberg model at a certain magnetization.
+The lowest-energy state of Eq.
+(5) coincides with the ground state of the Heisenberg model under a magnetic field.
+Due to the equivalence with the antiferromagnetic Heisenberg model, $\braket{\hat{\Delta}^{+}_{j+r}\hat{\Delta}^{-}_{j}}$ in Eq.
+(5) must have the phase factor $(-1)^{r}$, explaining the sign alternation of the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the photodoped ladder.
 Figure 4: (a) Linear scale plot and (b) log-log plot of the rung-rung DH pairing correlation function with varying $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$.
@@
 Inset: Correspondence between the photodoped ladder and the 1D antiferromagnetic Heisenberg model.
-The blue and purple ellipses represent the rung spin-singlet and rung DH $\eta$-triplet, respectively, as shown in Fig.
+The blue and purple symbols represent the rung spin singlet and rung DH $\eta$ triplet, respectively, as shown in Fig.
 1(b).
-This effective model assumes $V_{\perp}+J_{\perp}>|2t_{\perp}|$ to set the spin-singlet and $\eta$-triplet states to the lowest-energy configurations in the independent rung limit [86].
-Under this condition, $K>0$, and therefore Eq.
-(5) is equivalent to the 1D antiferromagnetic Heisenberg model at a certain magnetization.
-The lowest-energy state of Eq.
-(5) coincides with the ground state of the Heisenberg model under a magnetic field.
-Due to the equivalence with the antiferromagnetic Heisenberg model, $\braket{\hat{\Delta}^{+}_{j+r}\hat{\Delta}^{-}_{j}}$ in Eq.
-(5) must have the phase factor $(-1)^{r}$, explaining the sign alternation of the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the photodoped ladder.
 Going beyond the above analytical perturbative arguments, we study the stability of the sign alternation in $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{i}\hat{\Delta}^{\mathrm{r}}_{j}}$ as a function of the $t_{\parallel},J_{\parallel},V_{\parallel}$ parameters of the full model $\hat{\mathcal{H}}$ numerically.
-Figure 4 shows the DH pairing correlations as $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$ increase linearly, departing from ^$\hat{\mathcal{H}}_{\rm LRA}$.
+Figure 4 shows the DH pairing correlations as $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$ increase linearly, departing from $\hat{\mathcal{H}}_{\rm LRA}$.
 As expected from the effective two-level model in Eq.
 (5), the DH pairing correlation is well developed when different rungs are only weakly connected ($p=0$).
-Although the absolute value of the correlation gradually decreases at larger $p$ as the parameters approach the original $\hat{\mathcal{H}}$, the sign alternation in the correlation function remains, confirming that ^$\hat{\mathcal{H}}_{\mathrm{min}}$ captures the essential features of the pairing correlations observed in the DH pairing phase.
+Although the absolute value of the correlation gradually decreases at larger $p$ as the parameters approach the original $\hat{\mathcal{H}}$, the sign alternation in the correlation function remains, confirming that $\hat{\mathcal{H}}_{\mathrm{min}}$ captures the essential features of the pairing correlations observed in the DH pairing phase.
 […]
-In contrast to the $\eta$-pairing phase that prominently appears in the $V\sim 0$ and large $n_{d}$ region, the DH pairing state can appear in the large $V$ and small $n_{d}$ region.
-Because strong photoexcitation to achieve high $n_{d}$ is not required and $V$ is nonzero in real materials, the DH pairing state is more experimentally accessible.
+The DH pairing state can appear in the small $n_{d}$ region, thus strong photoexcitation to achieve high $n_{d}$ is not required, and heating effects should be less destructive.
+The Coulomb interaction $V$ cooperatively contributes to the emergence of the DH pairing state; such an interaction is natively present in many correlated materials.
 For instance, ladder-type cuprates [100, 101, 102, 103] may serve as hosts for the DH pairing state.
-While the implementation of nearest-neighbor Coulomb interaction is a challenging ingredient for the realization with cold atoms, the first step towards this has been achieved for the Bose-Hubbard model [97].
+Having charted the DH pairing in a 2D setting, various other materials, such as cuprates with the 2D square structure, hold the potential to exhibit DH pairing.
+While the implementation of a nearest-neighbor Coulomb interaction is a challenging ingredient for the realization with cold atoms, the first step towards this has been achieved for the Bose-Hubbard model [104].
 Although photodoping is inherently a nonequilibrium phenomenon, we have approximately mapped a photodoped state to the lowest-energy state in the pseudoequilibrium condition.
 To further validate the emergence of the DH pairing phase and its lifetime, it would be necessary to demonstrate the time-dependent behavior of the DH pairing correlations in the optically driven Hubbard ladder using time-evolution methods [105, 106, 107, 108, 109].
+Having obtained a promising behavior on a small 2D cylinder, further studies of the $d$-wave nature of pairing should be performed on larger 2D lattices.
 These investigations remain as future challenges.
-Acknowledgements.
+Acknowledgments.
 We thank S.
@@
 Murakami, S.
-Nishimoto, and T.
+Nishimoto, K.
+Pradhan, and T.
 Sato for fruitful discussions.
 This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grants No.
-JP20H01849, No.
-JP24K06939, No.
+JP20H01849 (T.K.), No.
+JP24K06939 (T.K.), No.
 JP24H00191 (T.K.), No.
@@
 R.U.
-was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program”.
+was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program” and JST SPRING Grant No.
+JPMJSP2138.
 M.S.
-acknowledges the support from the UK EPSRC award under the Agreement EP/Y005090/1.
+acknowledges the support from the U.K.
+EPSRC award under Agreement No.
+EP/Y005090/1.
 Z.L.
-acknowledges the support by the QuantERA II JTC 2021 grant T-NiSQ by MVZI, the P1-0044 program of the Slovenian Research Agency, and ERC StG 2022 project DrumS, Grant Agreement 101077265.
+acknowledges the support by the QuantERA II JTC 2021 Grant T-NiSQ by MVZI, the P1-0044 program of the Slovenian Research Agency, and ERC StG 2022 project DrumS, Grant Agreement No.
+101077265.
 D.G.
@@
 Our calculations were performed using the ITensor library [110, 111].
+Data availability. The data that support the findings of this article are openly available [112].